# Jet Discrimination with Quantum Complete Graph Neural Network

### 08 Mar 2024 v1

@@ I Introduction
In this paper, we introduce the ”Quantum Complete Graph Neural Network (QCGNN)”, a variational quantum circuit model [26, 27, 28] specifically designed for learning complete graphs [29], showcasing a polynomial speedup over its classical counterparts by utilizing the property of quantum parallelism. The application of QCGNN is studied through jet discrimination of two-prong jets ($H\rightarrow b\bar{b}$) and three-prong jets (fully hadronic top jets) against light quark ($u,d,c,s$) jets, where we represent jets using complete graphs.
@@ II.2 Quantum Complete Graph Neural Network
In the noisy intermediate-scale quantum (NISQ) era [34] of quantum computers, the variational quantum circuit (VQC) [26, 27, 28] is one of the intuitive ways to realize quantum neural networks. Typically, a VQC can be formulated as

### 12 Mar 2024 v2

@@ I Introduction
In this paper, we introduce the ”Quantum Complete Graph Neural Network (QCGNN)”, a variational quantum algorithm [26, 27, 28] based model specifically designed for learning complete graphs [29], showcasing a polynomial speedup over its classical counterparts by utilizing the property of quantum parallelism. The application of QCGNN is studied through jet discrimination of two-prong jets ($H\rightarrow b\bar{b}$) and three-prong jets (fully hadronic top jets) against light quark ($u,d,c,s$) jets, where we represent jets using complete graphs.
@@ II.2 Quantum Complete Graph Neural Network
In the noisy intermediate-scale quantum (NISQ) era [34] of quantum computers, the variational quantum algorithm [26, 27, 28] is one of the intuitive ways to realize quantum neural networks, where the tunable part is built with a variational quantum circuit (VQC) and the parameters are updated through classical subroutines. Typically, a VQC can be formulated as